OpenStudy (anonymous):

what does it mean to refer to a 20 inch tv set or a 25 in tv set? such units refer to the diagonalof the screen. a 25 in tv set also has a width of 2o inches what is its height in inches

OpenStudy (anonymous):

Hint: It's a right triangle. Where 25 is the hypotenuse and 20 is the vertical length.

OpenStudy (anonymous):

i was asking how to figure it out but i need to see how its worked

OpenStudy (anonymous):

Sure. Do you know the formula of a right triangle? (hint: do you know pythagorean's theorem?)

OpenStudy (anonymous):

not sure this is a new chapter we are learning

OpenStudy (anonymous):

did you hear about a squared plus b squared equals c squared?

OpenStudy (anonymous):

thats the theorum

OpenStudy (anonymous):

Good! So what does each letter mean? the a,b,c?

OpenStudy (anonymous):

(also, it's five am here, so I've gotta head to bed soon, please try to answer as quickly as possible)

OpenStudy (anonymous):

are we not looking for c in the problem

OpenStudy (anonymous):

I'm afraid not. A and B are 'legs', and C is the hypotenuse.

OpenStudy (anonymous):

What are we missing? One of the two legs or the hypotenuse/

OpenStudy (anonymous):

i would say the hypotenuse

OpenStudy (anonymous):

Well, just to double check: do you know what a hypotenuse is?

OpenStudy (anonymous):

the longest side of a right angle and the prolem is asking for the hieght

OpenStudy (anonymous):

Perfect. And the longest side is opposite of the right angle (you'll have to trust me on this if you haven't learned this yet).

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So let's imagine this: your right angle is the bottom right corner of your screen. And our diagnol line will run from the top right (where the x on your screen is) to your start/windows button.

OpenStudy (anonymous):

So far so good? so we have a long line running in this direction /

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sweet. So we know our diagnal is 25 and our width (where your start menu and clock are) is 20.

OpenStudy (anonymous):

So, plug these numbers in for me into our formula.

OpenStudy (anonymous):

(and also show me what your formula looks like after plugging it in)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

am i getting ahead or wouldnt both the height and width be squared\[25^{2} +20^{2 }\]

OpenStudy (anonymous):

kinda.

OpenStudy (anonymous):

c is our hypotenuse, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and do we have a value for our hypotenuse? (remember the rule about hypotenuse being opposite of our right angle)

OpenStudy (anonymous):

at the moment its not known

OpenStudy (anonymous):

is it really?

OpenStudy (anonymous):

its what we are trying to find

OpenStudy (anonymous):

Remember that imaginary diagnal I had you draw through your monitor?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

was that not a hypotenuse?

OpenStudy (anonymous):

i see your point

OpenStudy (anonymous):

So how much is our hypotenuse, then? :P

OpenStudy (anonymous):

Battery will die soon, so I'll go ahead and make an exception to my usual style of teaching and give you the formula:

OpenStudy (anonymous):

25

OpenStudy (anonymous):

\[20^2 + b^2 = 25^2\]

OpenStudy (anonymous):

400 + b squared = 625

OpenStudy (anonymous):

alright, gotta go. Subtract 400 from both sides and find out what squared = the new number. Good luck.

OpenStudy (anonymous):

225

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

would it be 225 inches?

OpenStudy (anonymous):

remember: B^2 (b squared) = 225 So how do you get B by itself?